OpenStudy (anonymous):

Match each pair of equivalent ratios. 1)10 to 25 2)6 to 8 3)13/20 4) 15:40 ______________ A) 65% B)3/4 C) 2:5 D) 6 to 16

OpenStudy (anonymous):

oh boiii lol

OpenStudy (anonymous):

This is sooo hard XD

OpenStudy (anonymous):

Help me.

OpenStudy (anonymous):

I'm not kidding it's hard for me.

OpenStudy (mathmate):

Convert them all to decimals, and then compare. Example: 10 to 25 => 10/25=0.4 6 to 8 = 6/8 = 0.75 etc. etc.

OpenStudy (anonymous):

And then see which ones have equivalent decimals? o.o

OpenStudy (anonymous):

pretty much! yep

OpenStudy (anonymous):

THANKS:) I'll be in chat soon after I do this.

OpenStudy (anonymous):

Thanks for the ones who helped;)